OpenStudy (anonymous):

Solve 3x^2 = –30x + 24. Its Quadratics, can someone work it out for me please? :)

Parth (parthkohli):

Hey! Welcome to OpenStudy! You you rearrange the whole thing to get the standard form \(3x^2 + 30x - 24 = 0\) We can factor out \(3\) and get \(3(x^2 + 10x - 8) = 0\) Divide both sides by \(3\) --- \(x^2 + 10x - 8 = 0\)

Parth (parthkohli):

Now solve that using the quadratic formula.

Parth (parthkohli):

\( \color{Black}{\Rightarrow x = \Large {-b \pm \sqrt{b^2 - 4ac} \over 2a}}\)

Parth (parthkohli):

@murkersona Are you there?

OpenStudy (anonymous):

I am here, I am writing it down to be able to remember it :)

OpenStudy (karatechopper):

good job parth!!